OpenStudy (anonymous):

Use a graph of the sine function to find the value of q for which sin Ө=0......any help?

OpenStudy (anonymous):

one thing that would be good to memorize is that \(\sin(0)=0\)

OpenStudy (anonymous):

I know that, but the graph is what I don't know how to draw :l

OpenStudy (anonymous):

here is the graph, you can see that sine is zero at integer multiples of \(\pi\)

OpenStudy (anonymous):

i mean "here is the graph" http://www.wolframalpha.com/input/?i=y+%3Dsin%28x%29

OpenStudy (anonymous):

it is periodic, with period \(2\pi\) it goes from 0 to 1 back to 0 then to -1 and back up to 0 in an interval of length \(2\pi\)

OpenStudy (anonymous):

thank you c:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Okai can I ask you one more question? How do I get Ө=xk